A car travels 300 km in the same time that a train travels 200 km. The speed of the car is 20 km per hour more than the speed of the train. Find the speed of the car and of the train.
Answer by lwsshak3(11628)

A car travels 300 km in the same time that a train travels 200 km. The speed of the car is 20 km per hour more than the speed of the train. Find the speed of the car and of the train.
let x=speed of the train
x+20=speed of the car
travel time=distance/speed
300/(x+20)=200/x
lcd:x(x+20)
300x=200(x+20)
300x=200x+4000
100x=4000
x=40
x+20=60
speed of the train=60 km/hr
speed of the car=40 km/hr